# Dynamical Schwinger Effect in Lead-Halide Perovskites

### v1

@@ Abstract
We observe strong photoluminescence of a lead-halide perovskite driven by deep sub-gap irradiation. Using the quasi-adiabatic Landau-Dykhne approach, we interpret this observation in terms of the dynamical Schwinger effect – tunneling ionization in strong fields. Further, the exponential sensitivity of the Schwinger effect to driving fields allows us to measure the local frozen-in fields in a nominally cubic MAPbBr3 single crystal at room temperature. Finally, we demonstrate an AC analogue of biasing in our system – the cooperation between two time-dependent fields simultaneously driving the sample. Our results (i) establish lead-halide perovskites as an excellent platform for simulating effects of strong fields on Dirac fields, (ii) contribute to the on-going discussion about inversion-breaking in MAPbBr3 single crystal, and (iii) pave the way for a mid-infrared light detection with lead-halide perovskites.
@@ I Introduction
One of the most important insights coming from the synthesis of Quantum Mechanics and Special Relativity was the realization that vacuum is not empty. A textbook demonstration of this remarkable fact would be an observation of an electric breakdown of an empty space. However, as was shown by Schwinger, this can only occur at extremely strong fields $E_{S}\approx 10^{20}$/m Schwinger (1951), which are currently beyond experimental reach. In view of the enormity of $E_{S}$ and its intuitive interpretation as the maximal static electric field attainable in principle, it is natural that the actual attempts to realize vacuum breakdown (“Schwinger Effect”, SE) focus on its dynamical versions where the vacuum pair creation is achieved under the influence of time-dependent fields Mourou et al. (2006). The experimental progress is motivated by a number of theoretical works that suggest to reduce the field amplitudes needed for the observation of SE by crafting profiles of the driving fields Popov (1971); Ringwald (2001); Schützhold et al. (2008). However, any theoretical proposal must rely on certain simplifying assumptions, which due to the inherently non-perturbative character of SE may be decisive to the outcome Gies and Klingmüller (2005); Torgrimsson et al. (2017). Given that current technology is not capable to test these ideas in practice, it is hard to tell at the moment how realistic will they ultimately turn out to be.
@@
In this context it appears desirable to be able to simulate the Schwinger effect and the various approaches to it in a more accessible setting. An obvious direction is a realization of field-induced tunneling ionization in a semiconductor with a Dirac-like band dispersion. It is straightforward to demonstrate then that the threshold field in this case would be $E_{s}\sim\Delta/ea\sim 10^{9}$/m Linder et al. (2018), which is readily achievable in modern laboratories (here $\Delta$, $a$ and $e$ are the band gap, unit lattice length and electron charge respectively). However, it turns out that the observation of this scaled-down version of SE also poses significant challenges. The main difficulty here lies in the fact that despite an abundance of far-reaching analogies Wehling et al. (2014), there is an important difference between the “true” Dirac field in vacuum and a Dirac material in that the former features relativistic invariance while the latter does not. The important consequence in question is that a free particle in vacuum can in principle be accelerated indefinitely, while a charge carrier in a semiconductor can produce secondary particles once it acquires a certain threshold kinetic energy comparable to the band-gap Jena (2022). Since the threshold for this process $E_{\mathrm{av}}\sim\Delta/e\lambda_{\mathrm{mfp}}$ is significantly lower than that for tunneling ionization (here $\lambda_{\mathrm{mfp}}>a$ is the mean free path of charged carriers in the material), the population of secondary charge carriers will tend to dominate over tunnel-ionized carriers, often ending up in a catastrophic avalanche breakdown of the material. As a result of these complications, the observations of tunneling ionization have been limited so far either to the cases where the relevant tunneling region is restricted to a narrow part of space to avoid avalanche breakdown (e.g. tunneling ionization in single atoms Raizer (1991) or tunnel diodes Esaki (1958) and tunnel junctions Giaever and Megerle (1961); Chen (2007) in electronics); or to extended systems with no gap Berdyugin et al. (2022). However, to establish a direct connection to the efforts in high energy physics, it is necessary to realize an analog Schwinger effect in an extended Dirac material with a finite bandgap.
@@
In this Article, we report the observation of dynamical tunneling ionization in a single-crystal sample of lead halide perovskite (LHP), which realizes a gapped Dirac system Jin et al. (2012). In order to suppress avalanche formation and emphasize tunneling, we note that the former needs a finite amount of time to build up while the latter occurs quasi-instantaneously. Therefore, we seek to induce ionization with alternating electric fields using frequencies that, on the one hand, correspond to photon energies significantly smaller than the band gap $\hbar\omega\ll\Delta$ (to ensure that we operate in the quasi-adiabatic regime); on the other hand, they should oscillate fast enough to avoid an avalanche breakdown ($\omega>eE/\sqrt{m\Delta}$, where $E$ and $m$ are the magnitude of the applied electric field and the band mass of the charge carrier, respectively). Additionally, we are careful to keep our irradiation intensity levels well below impact ionization thresholds for our pulse durations $\tau\approx 300$fs to suppress the population of secondary charge carriers Rethfeld et al. (2017). This is necessary to reveal the intrinsic tunneling processes in the material and also to be able to describe them using quasi-adiabatic methods.
@@
Figure 1: A) Schematic diagram of ionization across the energy gap $\Delta$ due to multiphoton process, and due to tunneling under the influence of different applied field amplitudes ($E_{1}$ and $E_{1}+E_{2}$). It can be seen that the magnitude of $E$ determines the width of the forbidden range, thus affecting the net tunneling rate exponentially (see the text for details); B) Tunneling under the influence of the of alternating fields $E_{1}(t)$ and $E_{2}(t)$ (yellow and red); due to exponential sensitivity of the tunneling rate (blue) on the magnitude of the field, the tunneling rate peaks at the maximum of the total field (magenta); consequently the total number of transitions (green) depends only on the sum of the amplitudes of applied time-dependent fields; C) PL spectrum of MAPbBr3 single crystal sample pumped by $\lambda=4\mu$m radiation; Inset: MAPbBr3 crystal with PL coming from the bulk of the sample; D) PL spectra under $\lambda=4\mu$m pumping, together with our theoretical prediction for weak fields (no fit parameters), and for intermediate electric fields (see Supplementary Material). Note that we normalize all curves to be unity at a given value of the field ($eaE_{AC}\simeq 0.272$eV), for convenience. E) An illustration of several most common channels of photo-carrier recombination with characteristic rate dependencies on the total number of charge carriers (bottom).
@@
Low irradiation intensities, however, imply that all measurements need to be performed in the regime where the concentration of tunnel-induced carriers is so low that standard transport measurements cannot be used for studying tunnel ionization. Therefore, our approach here is to infer the population of photo-excited carriers from the intensity of photoluminescence (PL) emitted upon their recombination with each other. The success of this approach crucially depends on the photoluminescent quantum efficiency of the parent material, namely the ratio of radiative recombinations to the total total number of excitations Braslavsky (2007). This consideration is the primary motivation for the choice of lead halide perovskites as the base for realizing the solid-state version of SE. Furthermore, effective behavior of lead halide perovskites renowned for excellent quantum optoelectronic properties Goetz et al. (2020) is described very accurately by a Dirac equation with a non-zero mass Becker et al. (2018); Volosniev et al. (2023a, b), strengthening the analogy between our findings and the physics of SE.
@@ II Observation of Dynamical Schwinger Effect in MAPbBr3
Lead-halide perovskites are known to feature strong sub-bandgap response in terms of two- and three-photon adsorption Zhou et al. (2021); Mei et al. (2022) whose analysis has been mainly focused around the photophysical properties of LHPs in the near-infrared regime as it can shed light onto surprising solar cell (photovoltaic) performance of LHPs Jena et al. (2019). At the same time, the response of the system to photons with lower frequency remains largely unexplored. In Fig. 1C we show the PL spectrum of MAPbBr3 illuminated with mid-infrared radiation ($\lambda\approx 4\mu$m). Given the wavelength of PL ($\lambda_{\mathrm{PL}}\approx 570nm$) which roughly corresponds to the bandgap energy in MAPbBr3 ($\Delta\approx 2.3$eV Saidaminov et al. (2015); Tilchin et al. (2016)) this process corresponds to a whopping frequency conversion factor of $\sim 10$ at the face value. Combined with the exponential sensitivity of this process to the intensity of the irradiating field in Fig. 1D, we are compelled to rationalize the observed PL under deep sub-gap irradiation as quasi-adiabatic tunnel ionization of electrons across the semiconductor band gap using the standard terminology of strong-field phenomena Kruchinin et al. (2018) (similar behavior was observed in CsPbBr3, which indicates the general character of the phenomenon; see Supplementary Material). This observation comprises the central finding of this work. In the following we will further substantiate this qualitative observation by a quantitative analysis of ionization in a periodically driven Dirac-like band structure.
@@ II.1 Quasi-adiabatic tunneling
Tunneling is the quintessential quantum effect whereby a system undergoes a transition that can be described semi-classically by a trajectory of which at least a part passes through a classically-forbidden region in the parameter space. As a result the rate of this transition is strongly suppressed, being exponentially sensitive to the width of the classically-forbidden region Landau and Lifshitz (1981), see Fig. 1A. For instance, in the case of static electric field $E$ the rate of tunneling ionization $W$ across the gap $\Delta$ can be shown to be
@@
with the characteristic cut-off field introduced in the previous sections $E_{s}\sim\Delta/ea$ Sauter (1931); Zener (1934); Schwinger (1951). A convenient way to extend this limiting expression to the problem of tunneling excitation of a semiconductor under the influence of time-dependent fields varying at the characteristic frequencies much smaller than those given by the bandgap is the so-called quasi-adiabatic Landau-Dykhne method Delone and Krainov (2000). Within this approach the ionization rate $W$ can be shown (see Supplementary Material) to be given as
@@
where $f(x)=\mathrm{arsinh}(x)-(\sinh(2\mathrm{arsinh}(x))-2\mathrm{arsinh}(x))/(8x^{2})$, and $\gamma_{K}$ is the so-called Keldysh parameter of the problem whose numeric value depends on the strength and the frequency of the driving field as well as on the details of the band structure of the material in question Keldysh (1965). In the specific case of lead-bromide perovskites, density functional theory calculations provide $\gamma_{K}\simeq 0.85\omega/(eaE_{AC})$, where $E_{AC}$ is the strength of the electric field induced by the laser, and $a\simeq 0.586$nm is the lattice constant MAPbBr3 in cubic phase Becker et al. (2018).
The value of Keldysh parameter determines the degree of adiabaticity of the problem, the two limiting cases being pure adiabatic tunneling ($\gamma_{K}\to 0$) and multiphoton absorption ($\gamma_{K}\to\infty$) as illustrated in Fig.1A. In our experiment $\gamma_{K}\lesssim 1$ (see Fig. 1D), which puts us in the quasi-adiabatic tunneling regime, where the tunneling rate can be well approximated by $W\sim\exp\left[-{1.4\Delta}/{(eaE_{AC})}\right]$ exhibiting the exponential dependence on the external electric field characteristic of the non-perturbative tunneling processes such as (static) Schwinger effect [cf. Eq. (1)].
@@ II.2 Comparison to the experiment
In this section we confirm the tunneling nature of the ionization by deriving the functional dependence of the PL intensity, $I_{\mathrm{PL}}$, on the strength of the pumping mid-infrared (mid-IR) radiation. Naively, one might assume $I_{\mathrm{PL}}\propto W$, i.e., every single tunneling event results in a PL photon. However this is only true if: 1) there are no non-radiative recombination channels; 2) the photo-excited electron-hole pairs remain well-separated in space. Neither of these conditions are satisfied in lead-halide perovskites. First of all, despite the excellent photoelectronic properties of LHPs the non-radiative recombination processes in them cannot be ignored altogether. Similarly, the long diffusion lengths of photocarriers in LHP imply strong overlap between electron-hole pairs which gives rise to a almost exclusively bi-molecular character of the recombination processes Richter et al. (2016); Stranks (2017).
@@
Here, the coefficients $\{A_{i}\}$ describe the rates of different decay channels of $n(t)$ such as: defect-assisted ($A_{1}$), bi-molecular (both radiative- and non-radiative; $A_{2}$), and Auger-type ($A_{3}$) recombination processes followed by higher-order processes which should in principle be taken into account for sufficiently high densities of charge carriers, see Fig. 1E.
In the limit of weak pumping fields, $n(t)$ remains small at all times, implying that $n(t)=n(0)\exp\left(-A_{1}t\right)$ and $I_{\mathrm{PL}}\sim n(0)^{2}$ where the initial population $n(0)$ is proportional to the tunneling ionization rate $W$. As the pumping intensity grows, the omission of higher-order terms in Eq. (3) is no longer justified. For example, the inclusion of both mono- and bi-molecular channels leads to $I_{PL}\sim\alpha W-\mathrm{ln}(1+\alpha W)$, where $\alpha$ is a fitting parameter (see Supplementary Material). This expression in turn is accurate if the densities of charge carriers are of the order of $10^{16}-10^{17}$cm-3 (see Supplementary Material), which occurs for $eaE_{AC}\lesssim 0.3$eV, providing an estimate for a number of excited particles in the system. For higher excitation densities, the channels beyond bi-molecular recombination have to be included.

### v2

# Dynamical Schwinger effect and non-perturbative light detection in lead halide perovskites

@@ Abstract
Dielectric breakdown of physical vacuum (Schwinger effect) is the textbook demonstration of compatibility of Relativity and Quantum theory. Although, the observation of this effect in its original static formulation is practically unachievable, it has been shown that the requirements on field strengths can be significantly reduced for dynamical generalizations of Schwinger effect. Here, we report on observation of an analog dynamical Schwinger effect in gapped Dirac semiconductor lead-halide perovskite MAPbBr3. Specifically, we observe strong photoluminescence of a lead-halide perovskite driven by deep sub-gap irradiation, and use the quasi-adiabatic Landau-Dykhne approach to interpret our data in terms of the dynamical Schwinger effect. Further, the exponential sensitivity of the Schwinger effect to driving fields allows us to measure the local frozen-in fields in a nominally cubic single perovskite crystal at room temperature. Finally, we demonstrate an AC analogue of biasing in our system – the non-perturbative cooperation between two time-dependent fields simultaneously driving the sample. Our results establish lead-halide perovskites as an excellent platform for simulating effects of strong fields on Dirac fields. In addition, they contribute to the on-going discussion about inversion-breaking in MAPbBr3 single crystal and pave the way for a mid-infrared light detection with lead-halide perovskites.
@@ I Introduction
One of the most important insights coming from the synthesis of Quantum Mechanics and Special Relativity is the realization that vacuum is not empty. Instead, it is rather a fluctuating sea of virtual particle-antiparticle pairs, which can be in principle made real in the presence of external fields Schwinger (1951). However, this process is negligible unless the fields are of the order of $E_{S}\approx 10^{20}$/m at which point particle-antiparticle creation leads to the dielectric breakdown of vacuum. In view of the enormity of $E_{S}$ and its intuitive interpretation as the maximal static electric field attainable in principle, it is natural that the actual attempts to realize vacuum breakdown (“Schwinger Effect”, SE) focus on its dynamical versions where the vacuum pair creation is achieved under the influence of time-dependent fields Mourou et al. (2006). The experimental progress is motivated by a number of theoretical works that suggest to reduce the field amplitudes needed for the observation of SE by crafting profiles of the driving fields Popov (1971); Ringwald (2001); Schützhold et al. (2008). However, any theoretical proposal must rely on certain simplifying assumptions, which due to the inherently non-perturbative character of SE may be decisive to the outcome Gies and Klingmüller (2005); Torgrimsson et al. (2017). Given that current technology is not capable to test these ideas in practice, it is hard to tell at the moment how realistic will they ultimately turn out to be.
@@
In this context it appears desirable to be able to simulate the Schwinger effect and the various approaches to it in a more accessible setting. An obvious direction is a realization of field-induced tunneling ionization in a semiconductor with a Dirac-like band dispersion. It is straightforward to demonstrate then that the threshold field in this case would be $E_{s}\sim\Delta/ea\sim 10^{9}$/m Linder et al. (2018), which is readily achievable in modern laboratories (here $\Delta$, $a$ and $e$ are the band gap, unit lattice length and the elementary charge, respectively). However, it turns out that the observation of this scaled-down version of SE also poses significant challenges. The main difficulty here lies in the fact that despite an abundance of far-reaching analogies Wehling et al. (2014), there is an important difference between the “true” Dirac field in vacuum and a Dirac material in that the former features relativistic invariance while the latter does not. The important consequence in question is that a free particle in vacuum can in principle be accelerated indefinitely, while a charge carrier in a semiconductor can produce secondary particles once it acquires a certain threshold kinetic energy comparable to the band-gap Jena (2022). Since the threshold for this process $E_{\mathrm{av}}\sim\Delta/e\lambda_{\mathrm{mfp}}$ is significantly lower than that for tunneling ionization (here $\lambda_{\mathrm{mfp}}>a$ is the mean free path of charged carriers in the material), the population of secondary charge carriers will tend to dominate over tunnel-ionized carriers, often ending up in a catastrophic avalanche breakdown of the material. As a result of these complications, the observations of tunneling ionization have been limited so far either to the cases where the relevant tunneling region is restricted to a narrow part of space to avoid avalanche breakdown (e.g. tunneling ionization in single atoms Raizer (1991) or tunnel diodes Esaki (1958) and tunnel junctions Giaever and Megerle (1961); Chen (2007) in electronics); or to extended systems with no gap Berdyugin et al. (2022). However, to establish a direct connection to the efforts in high energy physics, it is necessary to realize an analog Schwinger effect in an extended Dirac material with a finite bandgap.
@@
In this Article, we report the observation of dynamical tunneling ionization in a single-crystal sample of lead halide perovskite (LHP), which realizes a gapped Dirac system Jin et al. (2012). In order to suppress avalanche formation and emphasize tunneling, we note that the former needs a finite amount of time to build up while the latter occurs quasi-instantaneously. Therefore, we seek to induce ionization with alternating electric fields using frequencies that, on the one hand, correspond to photon energies significantly smaller than the band gap $\hbar\omega\ll\Delta$ (to ensure that we operate in the quasi-adiabatic regime); on the other hand, they should oscillate fast enough to avoid an avalanche breakdown ($\omega\gtrsim eE/\sqrt{m\Delta}$, where $E$ and $m$ are the magnitude of the applied electric field and the band mass of the charge carrier, respectively). Additionally, we are careful to keep our irradiation intensity levels well below impact ionization thresholds for our pulse durations $\tau\approx 300$fs to suppress the population of secondary charge carriers Rethfeld et al. (2017). Operation with relatively small intensities also allows us to avoid ‘memory’ effects in our system, which are extrinsic to tunneling ionization (see Supplementary Material). This is necessary to reveal the tunneling processes inherent in the material and also to be able to describe them using quasi-adiabatic methods.
@@
Figure 1: A) Schematic diagram of ionization across the energy gap $\Delta$ due to multiphoton process, and due to tunneling under the influence of different applied field amplitudes ($E_{1}$ and $E_{1}+E_{2}$). It can be seen that the magnitude of $E$ determines the width of the forbidden range, thus affecting the net tunneling rate exponentially (see the text for details); B) Tunneling under the influence of the of alternating fields $E_{1}(t)$ and $E_{2}(t)$ (yellow and red); due to exponential sensitivity of the tunneling rate (blue) on the magnitude of the field, the tunneling rate peaks at the maximum of the total field (magenta); consequently the total number of transitions (green) depends only on the sum of the amplitudes of applied time-dependent fields; C) PL spectrum of MAPbBr3 single crystal sample pumped by $\lambda=4\mu$m radiation; Inset: MAPbBr3 crystal with PL coming from the bulk of the sample; D) PL spectra under $\lambda=4\mu$m pumping, together with our theoretical prediction for weak fields (no fit parameters), and for intermediate electric fields (see the text for details). Note that we normalize all curves to be unity at a given value of the field ($eaE_{AC}\simeq 0.272$eV), for convenience. E) An illustration of several most common channels of photo-carrier recombination with characteristic rate dependencies on the total number of charge carriers (bottom).
@@
Low irradiation intensities, however, imply that all measurements need to be performed in the regime where the concentration of tunnel-induced carriers is so low that standard transport measurements cannot be used for studying tunnel ionization. Therefore, our approach here is to infer the population of photo-excited carriers from the intensity of photoluminescence (PL) emitted upon their recombination with each other. The success of this approach crucially depends on two factors. First, the energy gap of the material should be free of defects to enhance the probability of tunneling directly into the conduction band. Second, the parent material should enjoy high photoluminescence quantum efficiency, namely the ratio of radiative recombinations to the total number of excitations Braslavsky (2007). These considerations are the primary motivation for the choice of lead halide perovskites as the base for realizing the solid-state version of SE Kirchartz et al. (2020). Furthermore, effective behavior of lead halide perovskites renowned for excellent quantum optoelectronic properties Goetz et al. (2020) is described very accurately by a Dirac equation with a non-zero mass Becker et al. (2018); Volosniev et al. (2023a, b), strengthening the analogy between our findings and the physics of SE.
@@ II Observation of Dynamical Schwinger Effect in MAPbBr3
Lead-halide perovskites are known to feature strong sub-bandgap response in terms of two- and three-photon adsorption Zhou et al. (2021); Mei et al. (2022) whose analysis has been mainly focused around the photophysical properties of LHPs in the near-infrared regime as it can shed light onto surprising solar cell (photovoltaic) performance of LHPs Jena et al. (2019). At the same time, the response of the system to photons with lower frequency remains largely unexplored. In Fig. 1C we show the PL spectrum of MAPbBr3 illuminated with mid-infrared radiation ($\lambda\approx 4\mu$m). Given the wavelength of PL ($\lambda_{\mathrm{PL}}\approx 570nm$) which roughly corresponds to the bandgap energy in MAPbBr3 ($\Delta\approx 2.3$eV Saidaminov et al. (2015); Tilchin et al. (2016)) this process corresponds to a whopping frequency conversion factor of $\sim 10$ at the face value. Combined with the exponential sensitivity of this process to the intensity of the irradiating field in Fig. 1D, we are compelled to rationalize the observed PL under deep sub-gap irradiation as quasi-adiabatic tunnel ionization of electrons across the semiconductor band gap using the standard terminology of strong-field phenomena Kruchinin et al. (2018) (similar behavior was observed in CsPbBr3, which indicates the general character of the phenomenon; see Supplementary Material and Ref. Zhang et al. (2024)). This observation comprises the central finding of this work. In the following we will further substantiate this qualitative observation by a quantitative analysis of ionization in a periodically driven Dirac-like band structure.
@@ II.1 Quasi-adiabatic tunneling
Tunneling is the quintessential quantum effect whereby a system undergoes a transition that can be described semi-classically by a trajectory of which at least a part passes through a classically-forbidden region in the parameter space. As a result the rate of this transition is strongly suppressed, being exponentially sensitive to the width of the classically-forbidden region Landau and Lifshitz (1981), see Fig. 1A. For instance, in a static electric field, $E$, the rate of tunneling ionization $W$ across the gap $\Delta$ can be shown to be
@@
with the characteristic cut-off field introduced above $E_{s}\sim\Delta/ea$ Sauter (1931); Zener (1934); Schwinger (1951). A convenient way to extend this limiting expression to the problem of tunneling excitation of a semiconductor under the influence of time-dependent fields varying at the characteristic frequencies much smaller than those given by the bandgap is the so-called quasi-adiabatic Landau-Dykhne method Delone and Krainov (2000). Within this approach the ionization rate $W$ can be shown (see Methods) to be given as
@@
where $f(x)=\mathrm{arsinh}(x)-(\sinh(2\mathrm{arsinh}(x))-2\mathrm{arsinh}(x))/(8x^{2})$, and $\gamma_{K}$ is the so-called Keldysh parameter of the problem whose numeric value depends on the strength and the frequency of the driving field as well as on the details of the band structure of the material in question Keldysh (1965). In the specific case of lead-bromide perovskites, density functional theory calculations provide $\gamma_{K}\simeq 0.85\hbar\omega/(eaE_{AC})$, where $E_{AC}$ is the strength of the electric field induced by the laser, and $a\simeq 0.586$nm is the lattice constant MAPbBr3 in cubic phase Becker et al. (2018).
The value of the Keldysh parameter determines the degree of adiabaticity of the problem, the two limiting cases being pure adiabatic tunneling  [^1] Note that pure adiabatic tunneling competes with the avalanche breakdown. Indeed, the condition $\omega\gtrsim eE_{AC}/\sqrt{m\Delta}\simeq eaE_{AC}/\hbar$ is clearly violated in the limit $\gamma_{K}\to 0$. Here, we estimate $eE_{AC}/\sqrt{m\Delta}$ using parameters of MAPbBr3. ($\gamma_{K}\to 0$) and multiphoton absorption ($\gamma_{K}\to\infty$) as illustrated in Fig. 1A. In our experiment $\gamma_{K}\lesssim 1$ (see Fig. 1D), which puts us in the quasi-adiabatic tunneling regime, where the tunneling rate can be well approximated by $W\sim\exp\left[-{1.4\Delta}/{(eaE_{AC})}\right]$ exhibiting the exponential dependence on the external electric field characteristic of the non-perturbative tunneling processes such as (static) Schwinger effect [cf. Eq. (1)].
@@ II.2 Comparison to the experiment
In this section we confirm the tunneling nature of the ionization by deriving the functional dependence of the PL intensity, $I_{\mathrm{PL}}$, on the strength of the pumping mid-infrared (mid-IR) radiation. Naively, one might assume $I_{\mathrm{PL}}\propto W$, i.e., every single tunneling event results in a PL photon. However this is only true if: 1) there are no non-radiative recombination channels; 2) the photo-excited electron-hole pairs remain well-separated in space; 3) the density of charge carriers is low. Neither of these conditions are satisfied in our experiment. First of all, despite the excellent photoelectronic properties of LHPs the non-radiative recombination processes in them cannot be ignored altogether. Similarly, the long diffusion lengths of photocarriers in LHPs imply strong overlap between electron-hole pairs, which gives rise to a bi-molecular character of the recombination processes Richter et al. (2016); Stranks (2017).
@@
Here, the coefficients $\{A_{i}\}$ describe the rates of different decay channels of $n(t)$ such as: defect-assisted ($A_{1}$), bi-molecular (both radiative- and non-radiative; $A_{2}$), and Auger-type ($A_{3}$) recombination processes followed by higher-order processes, which should, in principle, be taken into account for sufficiently high densities of charge carriers, see Fig. 1E.
In the limit of weak pumping fields, $n(t)$ remains small at all times, implying that $n(t)=n(0)\exp\left(-A_{1}t\right)$ and $I_{\mathrm{PL}}\sim n(0)^{2}$ where the initial population $n(0)$ is proportional to the tunneling ionization rate $W$. This result is accurate as long as $n_{0}\ll A_{1}/A_{2}$. Note that $A_{1}/A_{2}\simeq 3.5\times 10^{16}$cm-3 Richter et al. (2016). As the pumping intensity grows, the omission of higher-order terms in Eq. (3) is no longer justified. For example, the inclusion of both mono- and bi-molecular channels leads to $I_{PL}\sim\alpha W-\mathrm{ln}(1+\alpha W)$, where $\alpha$ is a fitting parameter (see Methods). This expression is accurate for $eaE_{AC}\lesssim 0.3$eV, providing an estimate for a number of excited particles in the system $10^{16}-10^{17}$cm-3 at this field. For higher excitation densities, the channels beyond bi-molecular recombination have to be included.